OpenStudy (anonymous):

what would be the next number in the pattern? -96, -48, -24, -12

OpenStudy (anonymous):

-6 the formula is divided by 2 every time. -96/2 = -48/2 = -24 etc.

OpenStudy (solomonzelman):

Yeah, you divide by 2 to get every next term. \[\color{red} { (-96) } \color{blue} { \div 2 }\color{red} { = -48 } \]\[\color{red} { (-48) } \color{blue} { \div 2 }\color{red} { = -24 } \]\[\color{red} { (-24) } \color{blue} { \div 2 }\color{red} { = -12 } \]\[\color{red} { (-12) } \color{blue} { \div 2 }~~~\color{green} { = -6 } \]